# Open Source Implementations of Numerical Algorithms for

### v4

@@ Abstract
The complete elliptic integral of the first kind (CEI-1) plays in a significant role in mathematics, physics and engineering. There is no simple formula for its computation, thus numerical algorithms are essential for coping with practical problems involved. The commercial implementations for the numerical solutions, such as the functions ellipticK and EllipticK provided by MATLAB and Mathematica respectively, are based on ${\mathcal{K}}_{\mathrm{cs}}(m)$ instead of the usual form $K(k)$ such that ${\mathcal{K}}_{\mathrm{cs}}(k^{2})=K(k)$ and $m=k^{2}$. It is necessary to develop open source implementations for the computation of the CEI-1 in order to avoid potential risks of using commercial software and possible limitations due to the unknown factors. In this paper, the infinite series method, arithmetic-geometric mean (AGM) method, Gauss-Chebyshev method and Gauss-Legendre methods are discussed in details with a top-down strategy. The four key algorithms for computing CEI-1 are designed, verified, validated and tested, which can be utilized in R& D and be reused properly. Numerical results show that our open source implementations based on $K(k)$ are equivalent to the commercial implementation based on ${\mathcal{K}}_{\mathrm{cs}}(m)$. The general algorithms for computing orthogonal polynomials developed are significant byproducts in the sense of STEM education and scientific computation.
Keywords: Complete elliptic integral of the first kind (CEI-1); Orthogonal polynomials; Verification-Validation-Testing (VVT); Infinte series; Arithmetic-geometric mean (AGM); Gauss numerical integration
@@ 1 Introduction
  • converting the CEI-1 into a finite sum with $12n-$ 1 order algebraic precision by taking $n$-th order orthogonal polynomials via Gauss numerical integration approach.
@@ 2.1 Infinite Series Method
$\left\{\begin{split}(2j-1)!!&=(2j-1)(2j-3)\cdots 3\cdot 1\\ (2j)!!&=(2j)(2j-2)\cdots 4\cdot 2\\ 0!!&=(-1)!!=1\end{split}\right.$ (12)
are the double factorial. Therefore, we can find the infinite series for the CEI-1, viz.,
@@ 2.2 AGM Method
$\left\{\begin{split}u_{i+1}&=\frac{u_{i}+v_{i}}{2},\quad i\geq 0\\ v_{i+1}&=\sqrt{u_{i}\cdot v_{i}},\quad i\geq 0\\ u_{0}&=a\\ v_{0}&=b\end{split}\right.$ (18)
@@
There are lots of formulae about the AGM and the CEI-1, we just list the following necessary ones according to our requirements:
@@
$K(k)=\frac{\pi/2}{\operatorname{AGM}(1+k,1-k)}=\frac{\pi/2}{\operatorname{AGM}(1,\sqrt{1-k^{2}})}$ (27)
@@ 2.3 Gauss-Chebyshev Method
According to Appendix A.1, for the abstract function
@@
of the Chebyshev type defined by the integral of ${\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}}$ with parameter $k$, we can computed it by the extended Gauss-Chebyshev formulae of numerical integration. Let ${x}^{\mathrm{gc}}_{i}$ denote the $i$-th root of $T_{n}(x)=\cos(n\arccos x)$, i.e. the $n$-th order Chebyshev polynomial of the first kind, and ${w}^{\mathrm{gc}}_{i}$ denote the $i$-th weight corresponding to ${x}^{\mathrm{gc}}_{i}$, then we have
@@ 2.4 Gauss-Legendre Method
defined by the integral of $f(x,k)$ with parameter $k$ such that $f(\cdot,k)\in C^{2n}[a,b]$ can be computed by the extended Gauss-Legendre formulae of numerical integration. Let ${x}^{\mathrm{gl}}_{i}$ denote the $i$-th root of the $n$-th order Legendre polynomial $P_{n}(x)$ and ${w}^{\mathrm{gl}}_{i}$ denote the $i$-th weight corresponding to ${x}^{\mathrm{gl}}_{i}$ such that
@@
Although there are several equivalent definitions for the Legendre polynomial, the best way for computing the Legendre polynomial $P_{n}(x)$ is the following iterative formulae
@@
This formulae implies that we can compute $P_{n}(x)$ one by one with an iterative process.
@@ 3.1 Infinite Series Algorithm
The procedures for calculating CEI-1 with infinite series method is shown in Figure 3, in which CoefFun is a procedure name (pointer to function or similar entity) used to calling the procedure of computing $c_{j}$.
@@
There is a practical technique for computing $c_{j}$. The double factorials $(2j)!!$! and $(2j-1)!!$! may be large numbers if $j>10$ and the overflow problem may happen in some computer language if the type of data is integer. However, we can cope with the the value of fraction safely instead of dealing with its denominator and numerator separately. Actually, we have
@@
It should be noted that the function $\texttt{CoefFun}:\mathbb{N}\to\mathbb{R}$ is used as an argument for the procedure InfiniteSeries in Algorithm 3.1, which implies that the InfiniteSeries is a high order function and the pointer to function will be involved for the implementation with different computer programming languages. We remark that the pointer to function is a concept in the C programming language and its counterparts in Python and MATLAB are lambda expression and function handle respectively.
@@ 3.2 AGM Algorithm
Let $\mathbf{x}={[x_{1},x_{2}]}^{\textsf{T}}\in\mathbb{R}^{{2}\times{1}}$, then the AGM method can be converted to a 2-dim fixed-point problem with the iterative formulae
@@
5:     for $i\in\left\{1,\cdots,n\right\}$ do
@@
The iterative process of AGM implies a special kind of iteration of fixed-point, which can be computed in a generic way with the help of while-do or do-while loop. The procedure FixedPointSolver, as shown in Algorithm 3.2, computes the $n$-dim fixed-point iteratively. Note that the pointers to functions, viz. Update and Dist should be assigned with concrete procedures according to the data type of $\mathbf{x}$.
@@ 3.5 Compute CEI-1 with the Gauss-Legendre based on the roots, weights and kernel function
1:Pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, order $n$, $n$-dim root vector ${\mathbf{x}}^{\mathrm{gl}}$, variable $a$, variable $b$, and variable $k$.
@@
4:     for $i\in\left\{1,\cdots,n\right\}$ do  //${v}^{\mathrm{gl}}_{f,i}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$;
@@
1:The variable $k\in[0,1)$, positive integer $n$ with default value $n=50$
@@ 3.6 Generating the Weights Specified by the Legendre polynomial and its Derivative
4:     for $i\in\left\{1,\cdots,n\right\}$ do
@@ 3.7 Solving the $n$ Roots of the Legendre Polynomial $P_{n}(x)$
Figure 8 demonstrates the system of solving the root vector of $P_{n}(x)$. Usually in the books about special mathematical functions [8, 9, 10], the different kinds of orthogonal polynomials, such as the Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$ of the first or second kind, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$, are treated separately, which hides the general iterative formulae (58). Once the $i$-th root ${x}^{\mathrm{gl}}_{i}$ is obtained, the corresponding weight ${w}^{\mathrm{gl}}_{i}$ can be determined by (39).
@@
  • designing the update procedure NewtonUpdate in Algorithm 3.7 by using $P_{n}(x)$ with the iteration formulae in Newton’s method;
@@
8:     for $i\in\left\{1,2,\cdots,n\right\}$ do
@@ 3.8 Generating the Legendre polynomials and their derivatives
The general form of four typical different orthogonal polynomials (such as Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$), denoted by $Y_{n}(x)$, can be specified by the following iterative formulae
@@
  • firstly, we design general procedures OrthoPolynom in Algorithm 3.8 and DerivOrthoPolynom in Algorithm 3.8 for generating $Y_{n}(x)$ and $Y^{\prime}_{n}(x)$ respectively with the iterative formulaes directly;
@@
 Algorithm 27 Compute the orthogonal polynomials $Y_{n}(x)$ according to the iterative formulae. (OrthoPolynom)
@@
13:     for $k\in\left\{2,\cdots,n\right\}$ do
@@
 Algorithm 28 Compute the derivative $Y^{\prime}_{n}(x)$ of the orthogonal polynomial $Y_{n}(x)$ according to the iterative formulae (DerivOrthoPolynom).
1:The variable $x$, order $n\in\mathbb{N}$, constants $c_{0},c_{1}$ and $c_{2}$, three pointers to function A, B and C.
@@
18:     for $k\in\left\{2,\cdots,n\right\}$ do
@@
 Algorithm 29 Compute the coefficient $A(n)$ in the iterative formulae for Legendre polynomial (AnLegendre).
@@
 Algorithm 30 Compute the coefficient $B(n)$ in the iterative formulae for Legendre polynomial (BnLegendre).
@@
 Algorithm 31 Compute the coefficient $C(n)$ in the iterative formulae for Legendre polynomial (CnLegendre).
@@ 4 Verification, Validation and Testing
The verification of the four methods for solving CEI-1 is simple if each modules involved have been verified and validated. In Appendix D, we illustrate some typical test cases and omit the details for testing the modules for constructing the four numerical solvers for CEI-1 since it is simple and straight forward. After that, what we need do is just to compare the values of $K(k)$ obtained by Algorithm 3.1, Algorithm 3.2, Algorithm 3.3 and Algorithm 3.5.
@@
  • Our four numerical algorithms lead to the same output and the correctness can be verified by the Kseries since it is based on the expansion of formulae (14) which is right definitely and can be handled manually.
  • The available funcitons ellipticK in MATLAB and EllipticK in Mathematica have consistent numerical solutions, which implies that the algorithms adopted may be the same although the solutions are not acceptable.
@@
Why do we use the implicit parameter $n$ instead of the explicit parameter $\varepsilon$ to control the precision of computation in the GC-CEI-1 and GL-CEI-1 methods? The answer lies in the difficulty of estimating the upper bounds of $\left|{f}_{\mathrm{gc}}^{(2n)}(\xi,k)\right|$ and $\left|{f}_{\mathrm{gl}}^{(2n)}(\xi,k)\right|$ for $\xi\in(-1,1)$ and $k\in[0,\sqrt{2}/2]$. Consequently, it is difficult for us to design algorithms to estimate the ${n}^{\mathrm{gl}}_{f}$ and ${n}^{\mathrm{gc}}_{f}$ with the precision $\varepsilon$ directly. This feature gives us a good example to demonstrate the advantages and disadvantages of the methods and algorithms developed. It is valuable in the sense of education although it is a pity in the sense of engineering. For the readers who cares the controllability of the iteration numbers $n$ according to the precision $\varepsilon$ in the MP problem, we suggest the Newton-Costes method in numerical integration.
@@ 5 Conclusions
Theoretical analysis shows that the correctness of the numerical solution can be verified by the analytic formulae in the form of infinite series. As alternative choices, we proposed four feasible numerical schemes and designed the Kseries, Kagm, Kintgc and Kintgl solvers for CEI-1 to obtain consistent numerical results based on infinite series method, AGM method, Gauss-Chebyshev & Gauss-Legendre numerical integrations. The key algorithms developed have been verified, validated and tested.
Although our emphasis is put on the numerical algorithms for the CEI-1, the methods adopted can also be used for the complete elliptic integrals of the second and third kinds. Furthermore, the key modules designed for the four solvers are designed for more general problems so as to satisfy various applications as more as possible. Particularly, the iterative algorithms for generating orthogonal polynomials with iterative formulae $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ and their derivatives can be taken to compute the Legendre, Chebyshev, Laguerre and Hermite polynomials which have wide applications in mathematics, physics and engineering.

### v5

@@ Abstract
The complete elliptic integral of the first kind (CEI-1) plays a significant role in mathematics, physics and engineering. There is no simple formula for its computation, thus numerical algorithms are essential for coping with the practical problems involved. The commercial implementations for the numerical solutions, such as the functions ellipticK and EllipticK provided by MATLAB and Mathematica respectively, are based on ${\mathcal{K}}_{\mathrm{cs}}(m)$ instead of the usual form $K(k)$ such that ${\mathcal{K}}_{\mathrm{cs}}(k^{2})=K(k)$ and $m=k^{2}$. It is necessary to develop open source implementations for the computation of the CEI-1 in order to avoid potential risks of using commercial software and possible limitations due to the unknown factors. In this paper, the infinite series method, arithmetic-geometric mean (AGM) method, Gauss-Chebyshev method and Gauss-Legendre methods are discussed in details with a top-down strategy. The four key algorithms for computing CEI-1 are designed, verified, validated and tested, which can be utilized in R& D and be reused properly. Numerical results show that our open source implementations based on $K(k)$ are equivalent to the commercial implementation based on ${\mathcal{K}}_{\mathrm{cs}}(m)$. The general algorithms for computing orthogonal polynomials developed are significant byproducts in the sense of STEM education and scientific computation.
Keywords: Complete elliptic integral of the first kind (CEI-1); STEM education; Algorithm design; Orthogonal polynomials; Verification-Validation-Testing (VVT); Infinte series; Arithmetic-geometric mean (AGM); Gauss numerical integration
@@ 1 Introduction
  • converting the CEI-1 into a finite sum with the $12n-$ 1 order algebraic precision by taking the $n$-th order orthogonal polynomials via Gauss numerical integration approach.
@@ 2.1 Infinite Series Method
$\left\{\begin{split}&(2j-1)!!=(2j-1)(2j-3)\cdots 3\cdot 1\\ &(2j)!!=(2j)(2j-2)\cdots 4\cdot 2\\ &0!!=(-1)!!=1\end{split}\right.$ (12)
are the fundamental results for the double factorial. Therefore, we can find the infinite series for the CEI-1, viz.,
@@ 2.2 AGM Method
$\left\{\begin{split}u_{i+1}&=\frac{u_{i}+v_{i}}{2},\quad i\geq 0\\ v_{i+1}&=\sqrt{u_{i}\cdot v_{i}},\quad i\geq 0\\ u_{0}&=a\geq 0\\ v_{0}&=b\geq 0\end{split}\right.$ (18)
@@
There are lots of formula about the AGM and the CEI-1, we just list the following necessary ones according to our requirements:
@@
$K(k)=\frac{\pi/2}{\operatorname{AGM}(1,\sqrt{1-k^{2}})}$ (27)
@@ 2.3 Gauss-Chebyshev Method
According to the Appendix A.1, for the abstract function
@@
of the Chebyshev type defined by the integral of ${\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}}$ with parameter $k$, we can computed it by the extended Gauss-Chebyshev formula of numerical integration. Let ${x}^{\mathrm{gc}}_{i}$ denote the $i$-th root of $T_{n}(x)=\cos(n\arccos x)$, i.e. the $n$-th order Chebyshev polynomial of the first kind, and ${w}^{\mathrm{gc}}_{i}$ denote the $i$-th weight corresponding to ${x}^{\mathrm{gc}}_{i}$, then we have
@@ 2.4 Gauss-Legendre Method
defined by the integral of $f(x,k)$ with parameter $k$ such that $f(\cdot,k)\in C^{2n}[a,b]$ can be computed by the extended Gauss-Legendre formula of numerical integration. Let ${x}^{\mathrm{gl}}_{i}$ denote the $i$-th root of the $n$-th order Legendre polynomial $P_{n}(x)$ and ${w}^{\mathrm{gl}}_{i}$ denote the $i$-th weight corresponding to ${x}^{\mathrm{gl}}_{i}$ such that
@@
Although there are several equivalent definitions for the Legendre polynomial, the best way for computing the Legendre polynomial $P_{n}(x)$ is the following iterative formula
@@
This formula implies that we can compute $P_{n}(x)$ one by one with an iterative process.
@@ 3.1 Infinite Series Algorithm
The procedures for calculating CEI-1 with infinite series method is shown in Figure 3, in which CoefFun is a procedure name (pointer to function or similar entity) used to calling the procedure for computing $c_{j}$.
@@
There is a practical technique for computing $c_{j}$. The double factorials $(2j)!!$! and $(2j-1)!!$! may be large numbers if $j>10$ and the overflow problem may happen in some computer language if the type of data is integer. However, we can cope with the the value of the fraction safely instead of dealing with its denominator and numerator separately. Actually, we have
@@
It should be noted that the function $\texttt{CoefFun}:\mathbb{N}\to\mathbb{R}$ is used as an argument for the procedure InfiniteSeries in Algorithm 3.1, which implies that the InfiniteSeries is a high order function and the pointer to function will be involved for the implementation with different computer programming languages. We remark that the pointer to function is a concept in the C programming language and its counterparts in Lisp/Python/Java/C++ and MATLAB/Octave are lambda expression and function handle respectively.
@@ 3.2 AGM Algorithm
Let $\mathbf{x}={[x_{1},x_{2}]}^{\textsf{T}}\in\mathbb{R}^{{2}\times{1}}$, then the AGM method can be converted to a 2-dim fixed-point problem with the iterative formula
@@
5:     for $i\in\left\langle 1,\cdots,n\right\rangle$ do
@@
The iterative process of AGM implies a special kind of iteration for the fixed-point, which can be computed in a generic way with the help of while-do or do-while loop. The procedure FixedPointSolver, as shown in Algorithm 3.2, computes the $n$-dim fixed-point iteratively. Note that the pointers to functions, viz. Update and Dist should be assigned with concrete procedures according to the data type of $\mathbf{x}$.
@@ 3.5 Compute CEI-1 with the Gauss-Legendre based on the roots, weights and kernel function
1:The pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, the order $n$, the $n$-dim root vector ${\mathbf{x}}^{\mathrm{gl}}$, the variable $a$, the variable $b$, and the variable $k$.
@@
4:     for $i\in\left\langle 1,\cdots,n\right\rangle$ do  //${v}^{\mathrm{gl}}_{f,i}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$;
@@
1:The variable $k\in[0,1)$, the positive integer $n$ with default value $n=50$
@@ 3.6 Generating the Weights Specified by the Legendre polynomial and its Derivative
4:     for $i\in\left\langle 1,\cdots,n\right\rangle$ do
@@ 3.7 Solving the $n$ Roots of the Legendre Polynomial $P_{n}(x)$
Figure 8 demonstrates the system of solving the root vector of $P_{n}(x)$. Usually in the books about special mathematical functions [8, 9, 10], the different kinds of orthogonal polynomials, such as the Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$ of the first or second kind, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$, are treated separately, which hides the general iterative formula (58). Once the $i$-th root ${x}^{\mathrm{gl}}_{i}$ is obtained, the corresponding weight ${w}^{\mathrm{gl}}_{i}$ can be determined by (39).
@@
  • designing the update procedure NewtonUpdate in Algorithm 3.7 by using $P_{n}(x)$ with the iteration formula in Newton’s method;
@@
8:     for $i\in\left\langle 1,2,\cdots,n\right\rangle$ do
@@ 3.8 Generating the Legendre polynomials and their derivatives
The general form of four typical different orthogonal polynomials (such as Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$), denoted by $Y_{n}(x)$, can be specified by the following iterative formula
@@
  • firstly, we design general procedures OrthoPolynom in Algorithm 3.8 and DerivOrthoPolynom in Algorithm 3.8 for generating $Y_{n}(x)$ and $Y^{\prime}_{n}(x)$ respectively with the iterative formulas directly;
@@
 Algorithm 27 Compute the orthogonal polynomials $Y_{n}(x)$ according to the iterative formula. (OrthoPolynom)
@@
13:     for $k\in\left\langle 2,\cdots,n\right\rangle$ do
@@
 Algorithm 28 Compute the derivative $Y^{\prime}_{n}(x)$ of the orthogonal polynomial $Y_{n}(x)$ according to the iterative formula (DerivOrthoPolynom).
1:The variable $x$, the order $n\in\mathbb{N}$, the constants $c_{0},c_{1}$ and $c_{2}$, the three pointers to function A, B and C.
@@
18:     for $k\in\left\langle 2,\cdots,n\right\rangle$ do
@@
 Algorithm 29 Compute the coefficient $A(n)$ in the iterative formula for Legendre polynomial (AnLegendre).
@@
 Algorithm 30 Compute the coefficient $B(n)$ in the iterative formula for Legendre polynomial (BnLegendre).
@@
 Algorithm 31 Compute the coefficient $C(n)$ in the iterative formula for Legendre polynomial (CnLegendre).
@@ 4 Verification, Validation and Testing
The verification of the four methods for solving CEI-1 is simple if all of the modules involved have been verified and validated. In the Appendix D, we illustrate some typical test cases and omit the details for testing the modules for constructing the four numerical solvers for CEI-1 since it is simple and straight forward. After that, what we need do is just to compare the values of $K(k)$ obtained by Algorithm 3.1, Algorithm 3.2, Algorithm 3.3 and Algorithm 3.5.
@@
  • Our four numerical algorithms lead to the same output and the correctness can be verified by the Kseries since it is based on the expansion of formula (14) which is right definitely and can be handled manually.
  • The available funcitons ellipticK in MATLAB and EllipticK in Mathematica have consistent numerical solutions, which implies that the algorithms adopted may be the same.
@@
Why do we use the implicit parameter $n$ instead of the explicit parameter $\varepsilon$ to control the precision of computation in the GC-CEI-1 and GL-CEI-1 methods? The answer lies in the difficulty of estimating the upper bounds of $\left|{f}_{\mathrm{gc}}^{(2n)}(\xi,k)\right|$ and $\left|{f}_{\mathrm{gl}}^{(2n)}(\xi,k)\right|$ for $\xi\in(-1,1)$ and $k\in[0,\sqrt{2}/2]$. Consequently, it is difficult for us to design algorithms to estimate the ${n}^{\mathrm{gl}}_{f}$ and ${n}^{\mathrm{gc}}_{f}$ with the precision $\varepsilon$ directly. This feature gives us a good example to demonstrate the advantages and disadvantages of the methods and algorithms developed. It is valuable in the sense of education although it is a pity in the sense of engineering.
@@ 5 Conclusions
Theoretical analysis shows that the correctness of the numerical solution can be verified by the analytic formula in the form of infinite series. As alternative choices, we proposed four feasible numerical schemes and designed the Kseries, Kagm, Kintgc and Kintgl solvers for CEI-1 to obtain consistent numerical results based on infinite series method, AGM method, Gauss-Chebyshev & Gauss-Legendre numerical integrations. The key algorithms developed have been verified, validated and tested.
Although our emphasis is put on the numerical algorithms for the CEI-1, the methods adopted can also be used for the complete elliptic integrals of the second and third kinds. Furthermore, the key modules for the four solvers are designed for more general problems so as to satisfy various applications as much as possible. Particularly, the iterative algorithms for generating orthogonal polynomials with iterative formula $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ and their derivatives can be taken to compute the Legendre, Chebyshev, Laguerre and Hermite polynomials which have wide applications in mathematics, physics and engineering.